### v1

# Coupling of magnetic and lattice collective excitations in the 2D van der Waals antiferromagnet FePS3

Abstract
We report a comprehensive magneto-optical study of FePS3 single crystals using temperature-dependent polarized transmission spectroscopy and Faraday rotation measurements. Our data reveal strong in-plane anisotropy of both phonon and magnon excitations. These anisotropies emerge distinctly below the Néel temperature ($T_{N}\approx 118\leavevmode\nobreak\ K$). A prominent magnon mode at 122  $cm^{-1}$ exhibits characteristic redshifting with temperature and field-induced splitting, confirming its magnetic origin. The optical conductivity, extracted via Drude–Lorentz modeling, displays polarization-sensitive spectral changes across the magnetic transition. The observation of Raman-active modes in the infrared spectra and the detection of multiple phonons exhibiting Faraday rotation provide compelling evidence of spin–phonon coupling and inversion symmetry breaking in the magnetic state. These findings underscore the intricate interplay between magnetic and lattice degrees of freedom in FePS3 and establish its relevance for exploring spin–lattice interactions in two-dimensional antiferromagnets.
I Introduction
In recent years, atomically thin crystals, from graphene to transition-metal dichalcogenides, have transformed condensed-matter research and inspired a wide range of device concepts [1, 2]. Beyond their celebrated electronic properties, an expanding family of two-dimensional (2D) compounds displays intrinsic magnetism, unlocking fresh opportunities for spintronic and information-processing technologies [3, 4]. Consequently, major efforts have focused on reliable exfoliation, scalable synthesis, and device integration of 2D magnets. Despite these advances, fundamental questions remain about how magnetic order couples with lattice dynamics in 2D antiferromagnets, and whether unique optical signatures of spin–lattice interactions can be harnessed for probing or controlling magnetic phases.
Two-dimensional magnets are also an unparalleled platform for testing the fundamental physics of spin models. They permit direct access to Ising [5], XY [6, 7], Kitaev, and Heisenberg-type physics within a single atomic layer. This allows one to observe, manipulate, and exploit fluctuations of the magnetic order, and to probe how external stimuli modify the ground state [8, 9, 10, 11]. Robust (anti)ferromagnetic order at the 2D limit can emerge in the presence of magnetic anisotropy, interlayer coupling, or engineered perturbations. In practice, the resulting order and its excitations are exquisitely sensitive to the spin dimensionality, crystal-field anisotropy, and external tuning parameters such as strain, doping, and defects.
Magnetic van der Waals materials, including metal phosphorus trichalcogenides (MPX3) [M=Fe, Co, Ni, Mn; X=S, Se], represent a broad class of compounds with diverse spin dimensionality models, exhibiting both ferromagnetic and antiferromagnetic states. For example, FePS3 ($T_{\mathrm{N}}\approx 120$ K), MnPS3 ($T_{\mathrm{N}}\approx 78$ K), and NiPS3 ($T_{\mathrm{N}}\approx 155$ K), all with monoclinic crystal structures (space group C2/m), display distinct magnetic behaviors [12, 13]. FePS3 adheres to an Ising model with a zigzag antiferromagnetic configuration, featuring in-plane ferromagnetic chains that are antiparallel along one of the in-plane crystal axes, while NiPS3 and MnPS3 are best described by a XY and Heisenberg model, respectively [12, 13]. MnPS3 shows a Néel-type antiferromagnetic order, where neighboring atoms have antiparallel spins, in contrast to the zig-zag antiferromagnetic structure in FePS3. Notably, FePS3 maintains its magnetic order at the monolayer limit [14, 15] and undergoes an insulator-to-metal transition under quasi-hydrostatic pressure [16]. Inelastic neutron scattering measurements of FePS3 have identified two magnon excitations with finite gaps of 15 meV and 40 meV at $q=0$ [17]. Moreover, the comparable magnitudes of spin-orbit coupling and crystal field energies in FePS3 favor the presence of low-energy magnetic excitations [18]. Yet, a comprehensive understanding of how these low-energy magnetic excitations interact with phonons, what their circular conductivity spectral weight is, and how they affect crystal symmetries in FePS3 is still lacking.
The magnetic properties of these vdW materials are strongly coupled to the charge and lattice degrees of freedom and can be studied effectively using far-infrared spectroscopy. The additional advantage of optical spectroscopy is that it is also susceptible to the monolayer limit. Therefore, the magnetic state can be easily studied and detected using optical probes, such as Kerr/Faraday angle rotation, compared to other magnetometry techniques [19, 20]. The advantage of infrared spectroscopy is the direct access to the spectral weight of the different collective modes in the magnetic phase and, as a result, to the investigation of their coupling and relations to the underlying physics of the magnetic state.
In this article, we present a comprehensive infrared spectroscopy investigation of the 2D antiferromagnetic semiconductor FePS3. Using polarized infrared transmission spectroscopy and Faraday rotation measurements, we aim to elucidate the temperature and magnetic field dependence of both phonon and magnon excitations. Our primary objectives are to probe anisotropic optical responses, detect potential inversion symmetry breaking, and clarify the interplay between spin and lattice degrees of freedom in FePS3. To address these objectives, we synthesized high-quality FePS3 crystals and performed detailed magneto-optical measurements as described below.
II Methods
II.1 Sample preparation
FePS3 crystals were grown by the Chemical Vapor Transport (CVT) method, like other magnetic van der Waals materials in the MPX3 group [21]. A stoichiometric amount in the ratio 1:1:3 of elements Fe (99.9%, Chempure), P (99.99%, Alfa Aesar), and S (99.999%, Fluka), corresponding to approximately 0.5 gram total mass, were inserted in a quartz tube inside a glove box, together with approximately 2 mg/cm3 of iodine as transport agent. The tube was subsequently evacuated down to  10-4 mbar, with intermediate Ar flushing, and sealed to a length of 16 cm. The sealed tube was inserted in a horizontal furnace, in a controlled temperature gradient with 680 C and 630 C for the hot and cold ends, respectively. At the end of 20 days the furnace was switched off and the crystals were extracted from the tube.
II.2 Optical spectroscopy
To measure the magneto-optical coefficients, we have used a Bruker Vertex 70v Fourier Transform Infra-Red (FTIR) spectrometer coupled to a Cryogenic magneto-optical cryostat, allowing transmission and reflection measurements as a function of temperature, magnetic field, and incoming/outgoing polarization. We have used Hg arc lamp and Globar (SiC) sources for the far-infrared (FIR) and mid-infrared (MIR) frequency ranges, respectively. Residual magnetic fields affect the arc lamp, limiting reliable measurements below 40 $cm^{-1}$. Therefore, we primarily present results obtained with the Globar source. To detect the transmitted signal intensity in our measurements, we have used several liquid-helium-cooled Si bolometers from Infra-Red Labs, covering the range from about 10 cm-1 to about 4000 cm-1. The magnetic field of the cryostat can be tuned from -7 T to +7 T and was aligned parallel (perpendicular) to the light propagation vector (sample surface normal) in the Faraday configuration. The temperature range of the cryostat sample holder is from 4 K to about 300 K, while the sample is mounted with silver paste to the sample holder and situated in the vacuum chamber as the superconducting magnet during the measurement. The cryostat is positioned on an external linear translation stage, allowing the measurement of a sample in transmission (reflection) configuration and a bare hole (mirror) reference at any desired temperature, magnetic field, and incoming/outgoing polarization.
For the polarization-dependent measurements, we have installed gold or KRS polarizers for the FIR or MIR frequency ranges, respectively, along the incoming and outgoing light beam paths. The incoming polarizer was used for the temperature and magnetic field-dependent measurement for each principal polarization axis without an outgoing analyzer. In the Faraday rotation angle measurements, we fix the analyzer and rotate the incoming polarizer. Since FePS3 has a monoclinic crystal structure and shows an anisotropic optical response in the antiferromagnetic state, the typical fast protocol procedure [22], taking into account only four measurements, i.e. combination of external magnetic field ±$\pm B$ and $A=P=\pm 45^{\circ}$, where $P$ and $A$ are the incoming polarizer and outgoing analyzer angles, respectively, is subject to issues of numerical error. Therefore, we have applied the full protocol to measure the Faraday rotation angle for each of the two principal polarizations used in this work. In this protocol, the analyzer angle is fixed to one of the principal angles while measurement of the transmission intensity is being taken at intervals of $15^{\circ}$ for a full $360^{\circ}$ polarizer rotation. We have conducted the full protocol at a temperature of 5 K, i.e., deep in the antiferromagnetic state, and at fields of $B=\pm 7\leavevmode\nobreak\ T$. In addition, we have applied the fast protocol to investigate the temperature and magnetic field evolution of the Faraday angle rotation in FePS3.
III Results
III.1 Infrared Transmission
In Figure 1, we show the transmission color map measured along the two principal axes. On one hand, the transmission spectrum at temperatures substantially higher than the Néel antiferromagnetic transition temperature, $T_{N}=118\leavevmode\nobreak\ K$, of  FePS3 shows no major difference between the two principal axes (See in App. VI). On the other hand, a clear change in the spectra can be seen below the phase transition. Previous attempts to fit the transmission data of the non-polarized spectrum and therefore to extract the optical conductivity have failed, indicating the importance of the polarization dependence of the transmission spectrum and the anisotropy of the monoclinic crystal of  FePS3 in the magnetic state.
Figure 1: Polarized transmission through a 26 $\mu$m FePS3 flake as a function of frequency and temperature for two principal polarization (upper panel: Polarization 1, and lower panel: Polarization 2). A pronounced spectral evolution occurs below $T_{N}=118$ K, including the emergence of new phonons and redshifting of the 122 cm-1 magnon mode.
The transmission spectra in Fig. 1 exhibit several distinct changes in the phonon modes below $T_{N}$. For the first principal axis, the phonons’ linewidth narrows noticeably, while for the second axis, some phonons show a clear blueshift in their central frequency. In the latter case, we also detect a split in the phonon mode around 150 $cm^{-1}$. Table 1 details the comparison of the modes calculated by Density Functional Theory alongside the experimental values extracted from the absorption lines in the transmission data.
Table 1: Comparison of theoretical and experimental (Pol 1 and Pol 2) vibrational modes in FePS3. All values are in cm-1.
Symmetry Theoretical Pol 1 Pol 2
$\mathrm{B}_{u}$(1) 75.8
$\mathrm{B}_{u}$(2) 101.9 102.1
107.7
$\mathrm{B}_{u}$(3) 119.4
$\mathrm{A}_{u}$(1) 129.7 128.5 133.1
$\mathrm{A}_{u}$(2) 139.0 140.8 140.8
146.7
$\mathrm{A}_{u}$(3) 153.8 153.1
160.3
$\mathrm{B}_{u}$(4) 161.7 163.5
$\mathrm{B}_{u}$(5) 167.6 165.2 165.9
$\mathrm{B}_{u}$(6) 174.5 175.7 174.5
$\mathrm{A}_{u}$(4) 178.0
$\mathrm{B}_{u}$(7) 179.3 184.7 184.8
$\mathrm{B}_{u}$(8) 202.5
$\mathrm{A}_{u}$(5) 202.7
$\mathrm{A}_{u}$(6) 207.5 207.4
$\mathrm{A}_{u}$(7) 221.6 215.4 215.1
$\mathrm{B}_{u}$(9) 230.3 228.0 227.7
$\mathrm{B}_{u}$(10) 231.9 232.2
$\mathrm{B}_{u}$(11) 240.1
$\mathrm{B}_{u}$(12) 254.6 257.4 260.1
$\mathrm{B}_{u}$(13) 350.2 375.6
$\mathrm{B}_{u}$(14) 415.1
$\mathrm{B}_{u}$(15) 531.4 531.9
$\mathrm{A}_{u}$(8) 538.9
$\mathrm{B}_{u}$(16) 541.7 546.1
$\mathrm{B}_{u}$(17) 550.3
$\mathrm{A}_{u}$(9) 556.7 556.3 555.0
Along with the drastic change of the overall transmission spectra below $T_{N}$, a new mode appears at about 122  $cm^{-1}$, which seems to harden as temperature is lowered. We will now focus on the spectral weight, temperature, magnetic field, and polarization-induced changes in the collective excitations spectrum, which are coupled to the magnetic order.
III.2 Optical Conductivity
The real part of the optical conductivity, $\sigma_{1}$, was extracted from the Drude-Lorenz fit of a slab in a transmission model to the transmission raw data in the range of 40 to 4000  $cm^{-1}$ and is shown in Figure 2. We show the result of our analysis, in which we also employed the Variable Dielectric Function (VDF) method [23] to capture better the details of the raw transmission data at the cryostat’s base temperature of 5 K for both principal polarizations. The same procedure was used at all measured temperatures from 150 K down to 5 K to investigate the temperature dependence of the optical conductivity.
Figure 2: Real part of the optical conductivity $\sigma_{1}$ at 5 K in zero magnetic field, extracted via Drude–Lorentz modeling of transmission data for the two principal polarizations.
The transmission data and the real part of the optical conductivity depict the anisotropy in the phonon spectra detected only from temperatures close to the antiferromagnetic phase transition down to low temperatures deep in the magnetic state. As can be seen, the low-energy spectrum is different for the two principal polarizations chosen in this work. For example, the strong phonon mode at 143 $cm^{-1}$ of the 1st polarization is not present, at least with the same spectral weight, in the 2nd polarization spectra. The same goes for the phonon modes at 180 $cm^{-1}$ and 260 $cm^{-1}$  which seem to be centered around different energies and thus attributed to different vibrational modes.
III.3 FePS3 Magnon Excitation
Despite the anisotropy seen in the phononic spectra, the mode at 122  $cm^{-1}$, identified already as the low-energy 15 meV magnon of FePS3 [24, 25, 17], is present in both polarizations with the same spectral weight and at the same frequency. This suggests that the magnon in FePS3 is isotropic compared to the other lattice vibrational modes. Notably, the magnon at 122 $cm^{-1}$ appears in both the Raman shift and the infrared spectra. This observation suggests inversion symmetry breaking in FePS3, a feature not previously reported in the literature [26].
We will now turn to the temperature and magnetic field dependence of the 122  $cm^{-1}$ magnon excitation. As was shown above, the magnon hardens as the temperature is lowered and follows the same temperature dependence for both principal polarizations. Figure 3 shows the magnon’s central frequency as obtained by the Drude-Lorenz fit to our transmission data for both principal axes. To quantitatively describe the temperature dependence of the magnon frequency, we employed a BCS-like order parameter model, as given by the following expression:
$\omega_{0}(T)=a\cdot\tanh\left[1.82\left(1.018\left(\frac{T_{N}}{T}-1\right)\right)^{0.5}\right]+c$ (1)
where $a$ and $c$ are free parameters of the fit (see caption of Fig. 3 for values) and $T_{N}=118\leavevmode\nobreak\ K$. Eq. 1 (See Eq. 5 in Ref. [27]) is the Bardeen-Cooper-Schrieffer (BCS) type temperature dependence for the superconducting order parameter gap $\Delta$ with critical temperature $T_{c}$, i.e. $\Delta(T)=\Delta(0)\tanh(2.2\sqrt{(T_{c}-T)/T})$ [27, 28, 29]. Here, we show that the magnon central frequency $\omega_{0}$ follows the temperature dependence of a magnetic order parameter with transition temperature $T_{N}$ and provides clear evidence that the 122 $cm^{-1}$ excitation originates at the antiferromagnetic state.
Figure 3: Central frequency of the 122 cm-1 magnon as a function of reduced temperature ($T_{N}/T$) at zero magnetic field. The red line shows a fit to a BCS-like order parameter expression. The inset compares both principal polarizations, indicating isotropic temperature dependence.
Another compelling evidence for the magnetic origin of the 122 $cm^{-1}$ excitation was shown in recent Raman spectroscopy measurements of FePS3 [30] where the excitation, previously considered as a Brillouin-zone folded phonon branch [31, 32], splits into two modes as a function of the magnetic field. Figure 4 shows the field-induced splitting of the 122 $cm^{-1}$ magnon in our transmission spectra, consistent with similar observations in recent optical spectroscopy studies [30, 33]. This splitting becomes clearly resolved at fields up to 7 T. The magnon splitting follows a field dependence with a gyromagnetic ratio of a free electron, and as shown in Figure LABEL:fig:Omega_B for our data in the IR spectra, consistent with previous Raman measurements [30]. However, this field dependence breaks down at higher magnetic fields due to the avoided crossing of two modes, the lower energy magnon branch and the Raman mode at 103  $cm^{-1}$ [33].
Figure 4: Polarized magneto-transmission color maps of FePS3 at 5 K for two principal polarizations, measured as a function of magnetic field (up to 7 T in 1 T steps). Circles mark the central frequencies of the split magnon modes, highlighting the field-induced splitting behavior.
Figure 5: Magnetic-field dependence of the energy splitting $\Delta\nu$ between the central frequencies of the 122 cm-1 magnon mode in FePS3, measured for the two principal polarizations (Pol 1 and Pol 2) at 5 K. The splitting $\Delta\nu$ is extracted from fits to the transmission spectra and plotted as a function of magnetic field $B$. The dashed line represents the expected linear dependence assuming a gyromagnetic ratio ($\gamma\approx 0.94$ cm-1/T), which holds up to $\sim 7$ T before deviations arise due to collective modes’ interactions [33].
III.4 Faraday Angle Rotation
In addition to the infra-red spectroscopy of the temperature and magnetic field dependence of the 122 $cm^{-1}$ magnon excitation, we show in Figure 6 the Faraday angle rotation of FePS3 at 5 K and fields of $B=\pm 7\leavevmode\nobreak\ T$ which was not reported before, to the best of our knowledge. The measurement was obtained using the full protocol for the two different principal axis used in this work and as was explained in the previous section [Sec. II.2].
Figure 6: Faraday rotation angle (top panels) and corresponding magneto-transmission spectra (bottom panels) of FePS3 measured at 5 K under ±$\pm 7$ T magnetic field for two FIR polarizer orientations: (a,c) Pol 1 and (b,d) Pol 2. Pronounced Faraday rotation features appear at the 122 $cm^{-1}$ magnon, showing symmetric splitting with field, as well as at several phonon modes, highlighting strong spin–phonon coupling in the antiferromagnetic state. Transmission below 1% is masked (gray) to avoid unreliable Faraday angle extraction.
The Faraday angle rotation is shown in Fig. 6 along with the transmission as measured by the parallel polarization configuration of the analyzer and polarizer. To avoid artifacts in regions with very low signal, we masked out spectral regions where the transmission coefficient dropped below 1%. This threshold lies above the sensitivity limit of our experimental setup, ensuring the reliability of the Faraday angle measurements. In this way, we can discuss several interesting features, starting from the collective excitations which are originated at the antiferromagnetic spin configuration, i.e. magnons, and following with the lattice collective modes which are coupled to the magnetic order.
[…]
Figure 7: Temperature evolution of the Faraday rotation angle in the 280–350  $cm^{-1}$ range of FePS3, measured at 7 T with FIR polarizer orientation Pol 2. A prominent feature near 320  $cm^{-1}$ emerges below the Néel temperature (TN = 118 K), increases in strength with decreasing temperature, and is attributed to a higher-energy magnon mode. This temperature-dependent behavior further supports the magnetic origin of the excitation.
It is important to note here that the 122 $cm^{-1}$ magnon frequency in our data corresponds to the 15 meV magnon energy detected at the commensurate $q=(0,1,0)$ momentum vector via inelastic neutron scattering measurements [17] with a momentum dispersion reaching about 17 meV at $q=(0,0.5,0)$. On the other hand, the magnon at about 40 meV is seen at the incommensurate momentum of, e.g., $q=(\pm 1,k,\mp 1/3)$ with a momentum dispersion reaching about 17 meV at e.g., $q=(\pm 2,k,\mp 2/3)$ (where $k$ is $0.5,1,1.5$). In general, optical spectroscopy is sensitive only to $q\approx 0$ collective excitations. Nevertheless, the frequencies at which we measure the Faraday angle rotation signal agree with the magnon energies of FePS3 as measured by inelastic neutron scattering. Moreover, it is possible that the difference in intensity between the signal at 122 $cm^{-1}$ (15 meV) and 320 $cm^{-1}$ (40 meV) is associated with the fact that the excitations are commensurate and incommensurate, respectively.
IV Discussion
IV.1 Crystal Symmetries in FePS3
Several vibrational modes of FePS3 exhibit a clear Faraday angle rotation, providing strong evidence for spin–phonon coupling. This observation is consistent with earlier reports [24]. Interestingly, the Raman-active phonon mode at 108 $cm^{-1}$ appears as a weak absorption feature in the infrared magneto-transmission data, where it develops a measurable Faraday rotation of approximately 25 mrad. The same mode is also showing an avoided crossing behavior with the 122 $cm^{-1}$ magnon at high magnetic fields [24].
The appearance of the Raman modes in the IR data suggests a breaking of crystal symmetry. Application of magnetic field naturally breaks time-reversal symmetry. It hence can lead to the mixing of odd and even collective modes, such as those seen in our magneto-transmission measurements but also at the high field magneto transmission data of Ref.  [34], particularly the avoided crossing of the magnon low energy excitation and the Raman mode in the IR data. However, there is no conventional interpretation why this specific Raman mode will also develop a Faraday angle rotation.
Two additional modes show a different behavior, which also cannot be attributed to magnetic field-induced time-reversal symmetry (TRS) breaking. The first mode at 133 $cm^{-1}$ is an infrared active mode but develops a Faraday angle rotation of about 25 mRad. Therefore, this mode already shows a clear spin-phonon coupling through the Faraday angle rotation at low magnetic fields. The second mode at 166 $cm^{-1}$ is also intriguing, as it is a Raman-active mode that appears in the infrared transmission data even at zero magnetic field, but only below the antiferromagnetic transition temperature of FePS3 [30]. This mode also develops a Faraday rotation angle of approximately 50 mrad at 7 T, as shown in Fig. 6.
Figure 8 shows a color map of the polarized mid-infrared transmission data in our FePS3 single crystals and another intriguing spectral feature which is centered around 900 $cm^{-1}$. A wide absorption, and also a wide peak in the optical conductivity, can be seen mainly in the Pol2 data and below the AFM phase transition. Similarly, a peak in the Raman spectroscopy data of our samples is observed below the AFM phase transition and around a Raman shift of about 900 $cm^{-1}$.
Figure 8: Color map of polarized mid-infrared transmission at zero field as a function of temperature.
Figure 9: 2D color map of the Raman Intensity in zero field as a function of temperature.
Therefore, the appearance of the 122 $cm^{-1}$magnon, the 166 $cm^{-1}$aman-active phonon, and the mid-infrared excitation at 900 $cm^{-1}$, all visible in both infrared and Raman spectra at zero field, suggests that inversion symmetry is broken in FePS3. This claim is in line with recent experimental indications [35] and in contrast to second harmonic generation measurements [26], which argue that FePS3 preserves inversion symmetry in the magnetic state.
IV.2 Dichroism of the Optical Circular Conductivity
Since we can quantitatively obtain the absolute transmission, the complex conductivity, and the Faraday angle of FePS3 in our measurements, it is, therefore, possible to extract from our data the circular optical conductivity of the split magnon mode. The two field-separated collective modes can be associated with the right- and left-handed circular polarizations of the two opposite spin sublattices in the antiferromagnetic state  [36]. At zero field, there is no difference between one spin population and the other, and the magnon has a single eigen energy. Hence, the two circular polarizations combine into one linearly polarized mode at 122 $cm^{-1}$. It is important to note here that our measurements show no difference between the magnon mode at zero field and the two principal polarizations used in this work. We attribute it to the fact that FePS3 is considered to have an Ising-like (1D) spin Hamiltonian and that the spins are directed along the $c$-axis. Hence, there is no considerable effect on different polarizations along the crystal’s $ab$ plane. However, with the application of a magnetic field, the modes split into low and high energy modes with a corresponding circular optical conductivity $\sigma_{+}$ and $\sigma_{-}$.
We follow the procedure described in Ref. [22] however, due to the anisotropy in the magnetic state, the in-plane optical conductivity tensor should be described as follows:
$\sigma=\begin{pmatrix}\sigma_{xx}&\sigma_{xy}\\ -\sigma_{xy}&\sigma_{yy}\end{pmatrix}$ (2)
where the in-plane anisotropy of the FePS3 optical conductivity is given by different dynamic conductivity functions $\sigma_{xx}$ and $\sigma_{yy}$. The latter terms capture the ordinary absorption and the birefringence of the sample, while the diagonal term, $\sigma_{xy}$, is associated with the magneto-optical effects and the optical Hall conductivity. Plugging in the eigenvectors of the Right- and Left-handed circular polarized light $e_{\pm}=1/\sqrt{2}\left(1,\pm i\right)$ results in the circular optical conductivity:
$\sigma_{\pm}(\omega)=\frac{\sigma_{xx}(\omega)+\sigma_{yy}(\omega)}{2}\pm i\sigma_{xy}(\omega)$ (3)
for the Right-handed ($\sigma_{+}$) and Left-handed ($\sigma_{-}$) optical circular conductivity. It is important to note here that we work under the assumption that the monoclinic tilt is non-negligible and therefore take the orthorhombic approximation because the optic axes in our measurement stay nearly in-plane. We also note that $\sigma_{xy}$ should be independent of the actual axis in which we have resolved the Faraday angle rotation. Moreover, while linear birefringence mixes the two circular eigenmodes, so they are not perfectly circular, it does not affect the Faraday angle rotation since it is not related to the off-diagonal terms and cancels out in the cross-polarization measurement. However, the values of $\sigma_{xy}$ can be different for the measurement and analysis of the Faraday angle rotation in Pol1 vs Pol2 due to systematic errors, e.g. residual cross polarization signal, and/or crystal misalignment. Our results show that throughout most of the spectral range, including the lower energy magnon mode and phonons, there is a perfect agreement between the two sets of data with respect to $\sigma_{xy}$ obtained from the two in-plane polarizations. Some discrepancy between the values can be detected around the upper energy magnon mode (see App. VII), which shows that it is local in frequency and probably not related to systematic errors or any sample tilt. Therefore, we have averaged the two $\sigma_{xy}$ outputs in order to obtain the optical circular conductivity $\sigma_{\pm}$.
Figure 10 shows the circular optical conductivity obtained from the analysis of the Magneto-Optical Kerr/Faraday Effect [23] and by taking into account the in-plane optical conductivity tensor as described above. As expected from a collinear antiferromagnetic resonance, there is a clear dichroism between $\sigma_{+}$ and $\sigma_{-}$ under the application of a magnetic field seen at frequencies of about 116 $cm^{-1}$ and 129 $cm^{-1}$ for 7 T for the lower energy and upper energy, respectively, of the 122 $cm^{-1}$ zero-field magnon.
In fact, Fig.  10 shows that the lower mode is completely dichroic since the spectral weight of $\sigma_{-}$ in that excitation is the same magnitude as that of $\sigma_{1}$ at zero field while $\sigma_{+}$ shows a clear peak. On the other hand, the upper energy mode at around 129 $cm^{-1}$ does not show an ideal dichroic response with $\sigma_{+}$ being about factor of 1/3 from the value of $\sigma_{-}$. This response is not subject to the average of $\sigma_{xy}$ of the two polarizations and can be seen in both sets of data for the upper energy magnon mode.
Since this anomalous dichroism appears only in the upper energy magnon which is also close to an infra-red active phonon, we suggest that spin-phonon coupling is responsible for prohibiting optical transitions between the two circular energy states, thus giving a residual, but still significant, spectral weight to $\sigma_{+}$ which is supposed to be closer to $\sigma_{1}$ at zero field in that area (around $1\leavevmode\nobreak\ S$ $cm^{-1}$). Previous reports of coupling between the lower magnon mode and an optical phonon have been reported as an avoided crossing in the optical spectra [34]. Therefore, we suggest that the upper magnon mode is already developing an avoided transition due to spin-phonon coupling involving the upper energy magnon branch and the infrared active phonon at 133 $cm^{-1}$.
Figure 10: Circular optical conductivity $\sigma_{+}$ and $\sigma_{-}$ of FePS3 at 5 K and 7 T, obtained by averaging the off-diagonal conductivity $\sigma_{xy}$ extracted independently for Pol 1 and Pol 2 using Eq. 3.
A careful look into $\sigma_{\pm}$ exposes several other excitations which develop a dichroic response. Since the magnetic field is breaking the time-reversal symmetry, it is also possible to measure the circular dichroism of the Raman active mode at 108 $cm^{-1}$, which appears in our far infrared spectroscopy data and previous magneto-transmission measurements [34]. Figure 10 shows that the 108 $cm^{-1}$mode develops a Faraday angle rotation, which is above our measurement sensitivity. Another Raman active mode at 160 $cm^{-1}$ can be detected in the circular optical conductivity in both polarizations due to a possible inversion symmetry breaking and leakage of the even vibrational modes into the odd modes IR spectra. Since the 160 $cm^{-1}$mode is close to one of the IR active modes in our measurement, the Faraday angle of this mode is seen better in the second principal axis polarization shown in the right panel of Fig. 6. In both cases, our data allow us to extract the spectral weight of the circular conductivity of these modes.
Figure 11: Transmission spectra for two polarizations at (a) 150 K ($T>T_{N}$) and (b) 5 K ($T\ll T_{N}$).
V Conclusions
We have conducted a comprehensive infrared spectroscopic investigation of FePS3. By combining absolute transmission, optical conductivity, and Faraday rotation measurements, we explored its collective excitations. Our measurements reveal that the lattice vibrational modes exhibit marked anisotropy and field sensitivity in the antiferromagnetic state, providing direct evidence of strong spin–lattice coupling. A key feature is the 122 $cm^{-1}$ magnon mode, which appears in both infrared and Raman spectra. It exhibits isotropic temperature evolution and undergoes magnetic-field-induced splitting, consistent with antiferromagnetic order.
Significantly, several phonon modes display pronounced Faraday rotation, and certain Raman-active modes appear in the infrared spectra at zero field—features suggesting inversion symmetry breaking, likely driven by spin–phonon interactions. These observations challenge earlier reports suggesting that inversion symmetry is preserved in FePS3, highlighting the need for further studies to reconcile these findings. Moreover, our analysis of the circular optical conductivity reveals dichroism associated with magnon splitting. Anomalous behavior near certain phonon energies points to avoided crossings and magneto-elastic hybridization.
Our results reveal a rich magneto-optical landscape in FePS3 and demonstrate the critical role of spin–lattice interactions in shaping its collective excitation spectrum. Future experiments combining ultrafast optical probes or monolayer measurements could provide deeper insight into the dynamics of spin–lattice coupling and its potential for controlling magnetic phases in two-dimensional materials, and for their potential applications in spintronic and magneto-optical technologies.
Acknowledgments
This research was supported by the Israel Science Foundation (ISF) through project number 666/23. N.B. would like to acknowledge useful discussions with J. Teyssier, G. Blumberg, and M. Orlita.
VI Transmission Data
In Figure 11 we show a comparison of the transmission data measured for the two principal polarizations at 150 K, which is above the Néel temperature, TN=118 K, and at 5 K, which is deep in the antiferromagnetic state. At 150 K, there is no significant difference between the two spectra, except for the phonon at around 220 $cm^{-1}$, indicating isotropic transmission behavior in the paramagnetic phase and on the $a-b$ plane of the monoclinic structure of FePS3. However, as the temperature approaches TN, noticeable changes appear in the spectra, making it impossible to fit the data with a single isotropic model and therefore pointing to the onset of the magnetic order and increasing fluctuations of the magneto-elastic effects. The same spin fluctuations are also responsible for the change in the mode around 100 $cm^{-1}$seen in the Raman spectrum [30]. Below $T_{N}$, there is an abrupt change in the spectra up to about 250  $cm^{-1}$owing to the Brillouin zone folded phonons and appearance of the magnon mode, as was seen also previously in Raman spectroscopy.
VII Optical Circular Conductivity
Figure 12: Circular optical conductivities $\sigma_{+}$ and $\sigma_{-}$ reconstructed separately for each polarization (Pol 1 and Pol 2), using the combination of full-protocol Faraday rotation and absolute transmission spectra fitted using RefFit software.
Figure 13: Circular optical conductivities $\sigma_{+}$ and $\sigma_{-}$ reconstructed for each polarization (Pol 1 and Pol 2) using Eq. 3, where the Hall conductivity $\sigma_{xy}$ was independently derived for each polarization without averaging.
The main text considers an average of the dielectric constant in order to obtain the optical circular conductivity from the Faraday angle rotation and magneto transmission measurements in the two principal axes. Figures 12 and 13 provide complementary representations of the circular optical conductivity, prior to the averaging process, which help assess the consistency of the analysis presented in the main text (Fig. 10).
Figure 12 shows $\sigma_{+}$ and $\sigma_{-}$ reconstructed separately for the two principal polarizations (noted by the two anglea of the analyzer). In this case, the circular conductivities were obtained by combining the Faraday rotation measured using the full rotation protocol with the absolute transmission spectra, and fitting them directly using the MOKE + VDF procedure in RefFit software.
Figure 13 displays $\sigma_{+}$ and $\sigma_{-}$ for the two polarizations however now derived by using Eq. 3 where $\sigma_{xx}$ and $\sigma_{yy}$ were averaged. Here, the off-diagonal term $\sigma_{xy}$ was determined independently for each analyzer configuration without averaging.
Fig. 10 of the main text shows the result obtained by averaging $\sigma_{xy}$ from the two polarizations, providing a single optical circular conductivity spectrum for FePS3. Comparing Figs. 12 and 13 demonstrates that the overall important features of the circular dichroism that we noted in the main text, including the splitting of the magnon modes and the spectral weight asymmetry, are reproducible across all different analysis procedures. Minor differences between the reconstructions mainly reflect systematic effects such as residual cross-polarization signals and slight misalignment of the sample.