### v2

# Coupling of magnetic and lattice collective excitations in the 2D van der Waals antiferromagnet FePS3

Abstract
We combine polarized infrared magneto-transmission and Faraday angle rotation measurement to map the collective excitations of the van der Waals antiferromagnet FePS3. Below the Néel temperature ($T_{\mathrm{N}}\approx 118\leavevmode\nobreak\ K$), the phonon spectrum becomes strongly anisotropic, while a prominent excitation at $122$  $cm^{-1}$ ($15$ meV) hardens on cooling and splits linearly with magnetic field, identifying its magnetic origin. From absolute transmission and the Faraday angle, we reconstruct the circular optical conductivities, revealing a pronounced dichroism of the field-split excitations. The upper branch near $129$  $cm^{-1}$ shows an anomalously reduced dichroism consistent with spin–phonon hybridization with a nearby infrared phonon. Several phonons exhibit sizable Faraday rotation, and Raman-active modes appear in the infrared spectra; together with a broad mid-infrared band near $900$  $cm^{-1}$ emerging only below $T_{\mathrm{N}}$, these observations point to inversion-symmetry breaking in the magnetic state. A weak Faraday feature near $320$  $cm^{-1}$ ($\sim 40$ meV) further supports the presence of a higher-energy magnetic excitation. These findings underscore the intricate interplay between magnetic and lattice degrees of freedom in FePS3 and establish its relevance for exploring spin–lattice interactions in two-dimensional antiferromagnetic materials.
I Introduction
In recent years, atomically thin crystals, from graphene to transition-metal dichalcogenides, have transformed condensed-matter research and inspired a wide range of device concepts [1, 2]. Beyond their celebrated electronic properties, an expanding family of two-dimensional (2D) compounds displays intrinsic magnetism, unlocking fresh opportunities for spintronic and information-processing technologies [3, 4]. Consequently, major efforts have focused on reliable exfoliation, scalable synthesis, and device integration of 2D magnets. Despite these advances, fundamental questions remain about how magnetic order couples with lattice dynamics in 2D antiferromagnets, and whether unique optical signatures of spin–lattice interactions can be harnessed for probing or controlling magnetic phases.
Two-dimensional magnets are also an unparalleled platform for testing the fundamental physics of spin models. They permit direct access to Ising [5], XY [6, 7], Kitaev, and Heisenberg-type physics within a single atomic layer. This allows one to observe, manipulate, and exploit fluctuations of the magnetic order, and to probe how external stimuli such as temperature, field, strain, pressure modify the ground state [8, 9, 10, 11]. Robust (anti)ferromagnetic order at the 2D limit can emerge in the presence of magnetic anisotropy, interlayer coupling, or engineered perturbations. In practice, the resulting order and its excitations are exquisitely sensitive to the spin dimensionality, crystal-field anisotropy, and external tuning parameters such as strain, doping, and defects.
Magnetic van der Waals materials, including metal phosphorus trichalcogenides (MPX3) [M=Fe, Co, Ni, Mn; X=S, Se], represent a broad class of compounds with diverse spin dimensionality models, exhibiting both ferromagnetic and antiferromagnetic states. For example, FePS3 ($T_{\mathrm{N}}\approx 120$ K), MnPS3 ($T_{\mathrm{N}}\approx 78$ K), and NiPS3 ($T_{\mathrm{N}}\approx 155$ K), all with monoclinic crystal structures (space group C2/m), display distinct magnetic behaviors [12, 13]. FePS3 adheres to an Ising model with a zigzag antiferromagnetic configuration, featuring in-plane ferromagnetic chains that are antiparallel along one of the in-plane crystal axes, while NiPS3 and MnPS3 are best described by a XY and Heisenberg model, respectively [12, 13]. MnPS3 shows a Néel-type antiferromagnetic order, where neighboring atoms have antiparallel spins, in contrast to the zigzag antiferromagnetic structure in FePS3. Notably, FePS3 maintains its magnetic order at the monolayer limit [14, 15] and undergoes an insulator-to-metal transition under quasi-hydrostatic pressure [16]. Inelastic neutron scattering measurements of FePS3 have identified two excitations with finite gaps of 15 meV and 40 meV at $q=0$ [17]. Moreover, the comparable magnitudes of spin-orbit coupling and crystal field energies in FePS3 favor the presence of low-energy magnetic excitations [18]. Yet, a comprehensive understanding of how these low-energy magnetic excitations interact with phonons, what their circular conductivity spectral weight is, and how they affect crystal symmetries in FePS3 is still lacking.
The magnetic properties of these vdW materials are strongly coupled to the charge and lattice degrees of freedom and can be studied effectively using far-infrared spectroscopy. The advantage of optical spectroscopy is that it is also susceptible to the monolayer limit. Therefore, the magnetic state can be easily studied and detected using optical probes, such as Kerr/Faraday angle rotation, compared to other magnetometry techniques [19, 20]. The additional advantage of infrared spectroscopy is the direct access to the spectral weight of the different collective modes in the magnetic phase and, as a result, to the investigation of their coupling and relations to the underlying physics of the magnetic state.
In this article, we present a comprehensive infrared spectroscopy investigation of the 2D antiferromagnetic semiconductor FePS3. Using polarized infrared transmission spectroscopy and Faraday rotation measurements, we aim to elucidate the temperature and magnetic field dependence of the collective modes’ spectra. Our primary objectives are to probe anisotropic optical responses, detect potential inversion symmetry breaking, and clarify the interplay between spin and lattice degrees of freedom in FePS3. To address these objectives, we synthesized high-quality FePS3 crystals and performed detailed magneto-optical measurements as described below.
II Methods
II.1 Sample preparation
FePS3 crystals were grown by the Chemical Vapor Transport (CVT) method, like other magnetic van der Waals materials in the MPX3 group [21]. A stoichiometric amount in the ratio 1:1:3 of elements Fe (99.9%, Chempure), P (99.99%, Alfa Aesar), and S (99.999%, Fluka), corresponding to approximately 0.5 gram total mass, were inserted in a quartz tube inside a glove box, together with approximately 2 mg/cm3 of iodine as transport agent. The tube was subsequently evacuated down to  10-4 mbar, with intermediate Ar flushing, and sealed to a length of 16 cm. The sealed tube was inserted in a horizontal furnace, in a controlled temperature gradient with 680 C and 630 C for the hot and cold ends, respectively. At the end of 20 days the furnace was switched off and the crystals were extracted from the tube.
II.2 Infrared spectroscopy
To measure the magneto-optical coefficients, we have used a Bruker Vertex 70v Fourier Transform Infra-Red (FTIR) spectrometer coupled to a Cryogenic magneto-optical cryostat, allowing transmission measurements as a function of temperature, magnetic field, and incoming/outgoing polarization. We have used Hg arc lamp and Globar (SiC) sources for the far-infrared (FIR) and mid-infrared (MIR) frequency ranges, respectively. Residual magnetic fields affect the arc lamp, limiting reliable measurements below 40 $cm^{-1}$. Therefore, we primarily present results obtained with the Globar source. To detect the transmitted signal intensity in our measurements, we have used an IR Labs liquid-helium-cooled (and pumped) Si bolometer with cut-off filter for the range of 10 to 80  $cm^{-1}$, an IR Labs liquid-helium-cooled Si bolometer with cut-off filter for the range of 40 to 700  $cm^{-1}$, and an IR Labs liquid-helium-cooled Si bolometer with three different cut-off filters covering the range of 500 to 4000  $cm^{-1}$. The magnetic field of the cryostat can be tuned from -7 T to +7 T and was aligned parallel (perpendicular) to the light propagation vector (sample surface normal) in the Faraday configuration. The temperature range of the cryostat sample holder is from 4 K to about 300 K, while the sample is mounted with silver paste to the sample holder and situated in the vacuum chamber as the superconducting magnet during the measurement. The cryostat is positioned on an external linear translation stage, allowing the measurement of a sample in transmission (reflection) configuration and a bare hole (mirror) reference at any desired temperature, magnetic field, and incoming/outgoing polarization.
Three types of experiments were done in this work: (i) Polarization-dependent transmission measurements using gold or KRS-5 (thallium bromoiodide) polarizers for the FIR or MIR frequency ranges, respectively, along the incoming and outgoing light beam paths. Our preliminary attempts to measure and fit non-polarized transmission data and, as a result, to extract the optical conductivity have failed, indicating the importance of the polarization dependence of the transmission spectrum and the anisotropy of the monoclinic crystal of  FePS3 in the magnetic state. Therefore, we have used an incoming polarizer for the transmission measurements. As a first stage, we measured the FIR and MIR spectra as a function of polarizer angle deep in the magnetic state, at 5 K, and at zero magnetic field (See App. A). We have identified two distinct transmission spectra separated by a $90^{\circ}$ polarizer angle difference, to which we refer here as "Pol 1" and "Pol 2". We emphasize that these two principal polarizations were not associated with the aba-b plane crystal axis using x-ray diffraction apparatus and are only determined by the optical transmission spectra. Following this criterion, we have used these two principal polarizations for all the temperature and magnetic field-dependent transmission measurements without an outgoing analyzer. (ii) Faraday rotation angle measurements with a rotating polarizer at the incoming optical path and a fixed analyzer at the outgoing optical beam. The fast protocol was used to investigate the temperature and magnetic field evolution of the Faraday angle rotation in FePS3. Since FePS3 has a monoclinic crystal structure and shows an anisotropic optical response in the antiferromagnetic state, the typical fast protocol procedure [22], taking into account only four measurements, i.e. combination of external magnetic field $\pm B$ and $A=P=\pm 45^{\circ}$, where $P$ and $A$ are the incoming polarizer and outgoing analyzer angles, respectively, was subject to possible birefringence artifacts. Therefore, we have conducted (iii) full protocol Faraday rotation angle measurements for each of the two principal polarizations used in this work. In the full protocol, the analyzer angle is fixed to one of the principal angles while measurement of the transmission intensity is being taken at intervals of $15^{\circ}$ for a full $360^{\circ}$ polarizer rotation. We have conducted the full protocol at a temperature of 5 K, i.e., deep in the antiferromagnetic state, and at fields of $B=\pm 7\leavevmode\nobreak\ T$.
II.3 Raman microscopy
Several samples of FePS3 were measured in Raman spectroscopy using a Labram HR Evolution Horiba system combined with a CryoVac liquid helium flow cryostat. A 532-nm unpolarized laser was used as the excitation source and was focused on the sample (inside the cryostat) through a 63× Olympus objective, and the laser power was set to 0.4 mW throughout the experiment. The scattering light from the sample was collected by the same objective, passed through the analyzer, and sent to a Czerny-Turner spectrometer equipped with either a 600 or 1800 grooves per mm grating and was detected by a liquid nitrogen-cooled CCD-array. The spectra were obtained as a function of temperature from 300 K down to 10 K at temperature intervals of 10 K, in addition to a measurement at 5 K, which is the base temperature of the Cryostat in our system. The 1800 grooves per mm grating was used to measure the Raman shift spectrum in the range of about $\pm 600$ $cm^{-1}$ of Raman shift (i.e. Stokes and anti-Stokes transitions) around the elastic line and to resolve the frequency resolution of the different excitations, in particular the modes around 80 to 130 $cm^{-1}$appearing below the phase transition of FePS3. The 600 grooves per mm grating was used to measure the large Raman shift range of up to approximately 4000 $cm^{-1}$in order to resolve the excitations seen in our data and in previous reports on FePS3 around 1000 $cm^{-1}$ (e.g, Ref. [23]).
II.4 Phonon calculation
In order to obtain the phonon frequencies of FePS3  we have performed first-principles Density-functional theory (DFT) calculations with the Quantum ESPRESSO PWscf (Plane-Wave Self-Consistent Field) package v6.6 [24, 25, 26]. Exchange–correlation effects were treated within the PBE (Perdew-Burke-Ernzerhof) functional and the generalized-gradient approximation (GGA) [27]. Scalar-relativistic pseudopotentials (PPs) consistent with PBE were used: PAW (Projector-Augmented Wave) for Fe and US (Ultrasoft) for P and S [28]. The conventional unit cell and lattice parameters were adopted from the parameters of FePS3 in the Materials Project database [29, 30]. To capture the antiferromagnetic ground state, we used collinear spin polarization ($n_{\mathrm{spin}}=2$) with two Fe sublattices initialized antiparallel (Fe1/Fe2 starting magnetization $\pm 0.3125$), and small seed magnetizations on P and S (0.1). The converged self-consistent solution has essentially zero net moment with total magnetization $\approx 0\leavevmode\nobreak\ \mu_{\mathrm{B}}$ per cell (absolute magnetization $\sim 15\leavevmode\nobreak\ \mu_{\mathrm{B}}$). Wavefunctions and charge density were expanded with kinetic-energy cutoffs of 90 Ry and 1080 Ry, respectively. Metallic occupations were treated using Marzari–Vanderbilt cold smearing with a width of 0.0147 Ry ($\approx$ 0.20 eV) [31]. Brillouin-zone sampling employed a Monkhorst–Pack $6\times 4\times 3$ mesh [32]. Symmetry was enabled and four symmetry operations (including inversion, two with fractional translations) were detected. Electronic self-consistency used plain mixing with a mixing parameter $\beta=0.4$ and a target energy change $\mathrm{conv_{thr}}=8\times 10^{-9}$ Ry; in this calculation, convergence was reached in 20 SCF iterations. Calculations were carried out for 20 atoms per cell with four atomic species. Phonons were computed within density-functional perturbation theory (DFPT) as implemented in QE’s ph.x [33] and resulted in zone-center ($\Gamma$-point) phonon frequencies.
III Results
III.1 Infrared Transmission
In Figure 1, we show the transmission color map measured along the two principal axes. On one hand, the transmission spectrum at temperatures substantially higher than the Néel antiferromagnetic transition temperature, $T_{N}=118\leavevmode\nobreak\ K$, of  FePS3 shows no major difference between the two principal axes (See in App. A). On the other hand, a clear change in the spectra can be seen below the phase transition.
[…]
The transmission spectra in Fig. 1 exhibit several distinct changes in the phonon modes below $T_{N}$. For the first principal axis, the phonons’ linewidth narrows noticeably, while for the second axis, some phonons show a clear blueshift in their central frequency. In the latter case, we also detect a split in the phonon mode around 150 $cm^{-1}$. Table 1 details the comparison of the modes calculated by Density Functional Theory alongside the experimental values extracted from the absorption lines in the transmission data.
Table 1: Comparison of theoretical (a our work, b adapted from Ref. [34]) and experimental (Pol 1 and Pol 2) vibrational modes in FePS3. Modes labeled with are identified as Raman active modes in theory (e.g. Ref. [34]) and observed in the experimental data. Modes labeled with are identified with an additional symmetry in Ref. [34]. Modes labeled with c were also observed experimentally in Ref. [35]. All values are in cm-1.
Symmetry Th.a Th.b Pol 1 Pol 2
$\mathrm{B}_{u}$ 75.8
$\mathrm{B}_{u}$ 101.9 102.1
$\mathrm{A}_{g}$$\mathrm{B}_{g}$ 107.7
$\mathrm{B}_{u}$ 119.4
$\mathrm{A}_{u}$ 129.7 128.5 133.1
$\mathrm{A}_{u}$ 139.0 140.8c 140.8c
146.7
$\mathrm{A}_{u}$, $\mathrm{B}_{u}$ 153.8 150 153.1c
$\mathrm{A}_{g}$$\mathrm{B}_{g}$ 153
160.3c
$\mathrm{B}_{u}$ 161.7 158 163.5
$\mathrm{B}_{u}$ 167.6 165.2 165.9
$\mathrm{B}_{u}$ 174.5 175.7 174.5
$\mathrm{A}_{u}$ 178.0
$\mathrm{B}_{u}$ 179.3 184.7c 184.8c
$\mathrm{B}_{u}$ 202.5
$\mathrm{A}_{u}$ 202.7
$\mathrm{A}_{u}$ 207.5 207.4
$\mathrm{A}_{u}$, $\mathrm{B}_{u}$ 221.6 221 215.4c 215.1c
$\mathrm{B}_{u}$ 230.3 228.0 227.7
$\mathrm{B}_{u}$ 231.9 232.2
$\mathrm{B}_{u}$ 240.1
$\mathrm{A}_{u}$ 254.6 252
$\mathrm{B}_{u}$, $\mathrm{A}_{u}$ 254.6 258 257.4c
$\mathrm{B}_{u}$ 262 260.1
$\mathrm{B}_{u}$ 350.2 375.6
$\mathrm{B}_{u}$ 415.1 437
$\mathrm{B}_{u}$ 531.4 531.9
$\mathrm{A}_{u}$ 538.9
$\mathrm{B}_{u}$ 541.7 546.1
$\mathrm{B}_{u}$ 550.3
$\mathrm{A}_{u}$, $\mathrm{B}_{u}$ 556.7 570 556.3 555.0
Along with the drastic change of the overall transmission spectra below $T_{N}$, a new mode appears at about 122  $cm^{-1}$, or 15 meV, which seems to harden as the temperature is lowered. We will now focus on the spectral weight, temperature, magnetic field, and polarization-induced changes in the collective excitations spectrum, which are coupled to the magnetic order.
III.2 Optical Conductivity
The real part of the optical conductivity, $\sigma_{1}$, was extracted from the Drude-Lorentz fit of a plan-parallel slab transmission model to the transmission spectra measured in the range of 40 to 4000  $cm^{-1}$  and is shown in Figure 2. We show the result of our analysis, in which we also employed the Variable Dielectric Function (VDF) method [36] to capture better the details of the raw transmission data at the cryostat’s base temperature of 5 K for both principal polarizations. The same procedure was used at all measured temperatures from 150 K down to 5 K to investigate the temperature dependence of the optical conductivity.
Figure 2: Real part of the optical conductivity $\sigma_{1}$ at 5 K in zero magnetic field, extracted via Drude–Lorentz modeling of transmission data for the two principal polarizations.
The transmission data and the real part of the optical conductivity depict the anisotropy in the phonon spectra detected only from temperatures close to the antiferromagnetic phase transition down to low temperatures deep in the magnetic state. As can be seen, the low-energy spectrum is different for the two principal polarizations chosen in this work. For example, the strong phonon mode at 143 $cm^{-1}$ of the 1st polarization is not present, at least with the same spectral weight, in the 2nd polarization spectra. The same goes for the phonon modes at 180 $cm^{-1}$ and 260 $cm^{-1}$  which seem to be centered around different energies and thus attributed to different vibrational modes.
III.3 The 122  $cm^{-1}$ Excitation in FePS3
Despite the anisotropy seen in the phononic spectra, the mode at 122  $cm^{-1}$, identified previously as the 122  $cm^{-1}$ magnon of FePS3 [37, 38, 17], is present in both polarizations with the same spectral weight and at the same frequency. This suggests that the 122  $cm^{-1}$ excitation in FePS3 is isotropic compared to the other lattice vibrational modes. Notably, the excitation at 122  $cm^{-1}$ appears in both the Raman shift and the infrared spectra as shown in Fig. 3. This observation suggests inversion symmetry breaking in FePS3, a feature not previously reported in the literature [39].
Figure 3: Comparison of the Raman microscopy spectrum (red) with the transmission spectra for the two principal optical polarizations (black and blue) measured at 5 K and zero magnetic field. The comparison shows the 122 $cm^{-1}$ excitation in all spectra (The negligible slight discrepancy in the Raman shift is acceptable within our systematic error and can be attributed to the calibration of the elastic line in the Raman spectrometer). The figure shows several other excitations discussed in the text, namely the 133 $cm^{-1}$ and 166 $cm^{-1}$ IR modes and the 108 $cm^{-1}$ Raman mode.
We will now turn to the temperature and magnetic field dependence of the 122  $cm^{-1}$ excitation. As was shown above, this mode hardens as the temperature is lowered and follows the same temperature dependence for both principal polarizations. Figure 4 shows the mode’s central frequency as obtained by the Drude-Lorenz fit to our transmission data for both principal axes. Since the temperature dependence resembles that of an order parameter change, we employed the following expression to fit our data:
$\omega_{0}(T)=a\cdot\tanh\left[1.82\left(1.018\left(\frac{T_{N}}{T}-1\right)\right)^{0.5}\right]+c$ (1)
where $a$ and $c$ are free parameters of the fit (see caption of Fig. 4 for values) and $T_{N}=118\leavevmode\nobreak\ K$ (See Eq. 5 in Ref. [40] and relevant examples in Ref. [41, 42]). The mode’s central frequency $\omega_{0}$ follows the temperature dependence of a magnetic order parameter with transition temperature $T_{N}$ and provides clear evidence that the 122  $cm^{-1}$ excitation originates at the antiferromagnetic state.
Figure 4: Central frequency of the 122 cm-1 excitation as a function of reduced temperature ($T_{N}/T$) at zero magnetic field. The red line shows a fit to an order parameter temperature dependence (Eq. 1). The inset compares both principal polarizations, indicating isotropic temperature dependence.
Another compelling evidence for the magnetic origin of the 122  $cm^{-1}$ excitation was shown in recent Raman spectroscopy measurements of FePS3 [43] where the excitation, previously considered as a Brillouin-zone folded phonon branch [44, 45], splits into two modes as a function of the magnetic field. Figure 5 shows the field-induced splitting of the 122  $cm^{-1}$ mode in our transmission spectra, consistent with similar observations in recent optical spectroscopy studies [43, 46]. This splitting becomes clearly resolved at fields up to 7 T. The mode’s splitting follows a field dependence with a gyromagnetic ratio of a free electron, and as shown in Figure 6 for our data in the IR spectra, consistent with previous Raman measurements [43]. However, this field dependence breaks down at higher magnetic fields due to the avoided crossing of two modes, the lower energy mode’s branch and the Raman mode at 103  $cm^{-1}$ [46].
Figure 5: Polarized magneto-transmission color maps of FePS3 at 5 K for two principal polarizations, measured as a function of magnetic field (up to 7 T in 1 T steps). Circles mark the central frequencies of the split 122  $cm^{-1}$ mode, highlighting the field-induced splitting behavior.
Figure 6: Magnetic-field dependence of the energy splitting $\Delta\nu$ between the central frequencies of the 122 cm-1 mode in FePS3, measured for the two principal polarizations (Pol 1 and Pol 2) at 5 K. The splitting $\Delta\nu$ is extracted from fits to the transmission spectra and plotted as a function of magnetic field $B$. The dashed line represents the expected linear dependence assuming a gyromagnetic ratio ($\gamma\approx 0.94$ cm-1/T), which holds up to $\sim 7$ T before deviations arise due to collective modes’ interactions [46].
III.4 Faraday Angle Rotation
In addition to the infra-red spectroscopy of the temperature and magnetic field dependence of the 122 $cm^{-1}$ excitation, we show in Figure 7 the Faraday angle rotation of FePS3 at 5 K and fields of $B=\pm 7\leavevmode\nobreak\ T$ which was not reported before, to the best of our knowledge. The measurement was obtained using the full protocol for the two different principal axis used in this work and as was explained in the previous section [Sec. II.2].
Figure 7: Faraday rotation angle (top panels) and corresponding magneto-transmission spectra (bottom panels) of FePS3 measured at 5 K under $\pm 7$ T magnetic field for two FIR polarizer orientations: (a,c) Pol 1 and (b,d) Pol 2. Pronounced Faraday rotation features appear at the 122 $cm^{-1}$ excitation, showing symmetric splitting with field, as well as at several phonon modes, highlighting strong spin–phonon coupling in the antiferromagnetic state. Transmission below 1% is masked (gray) to avoid unreliable Faraday angle extraction.
The Faraday angle rotation is shown in Fig. 7 along with the transmission as measured by the parallel polarization configuration of the analyzer and polarizer. To avoid artifacts in regions with very low signal, we masked out spectral regions where the transmission coefficient dropped below 1%. This threshold lies above the sensitivity limit of our experimental setup, ensuring the reliability of the Faraday angle measurements. In this way, we can discuss several interesting features, starting from the collective excitations that originate at the antiferromagnetic state and then by the lattice collective modes that are coupled to the magnetic order.
The 122 $cm^{-1}$ excitation that splits into two modes shows a Faraday angle rotation of about 0.1 to 0.15 rad at 5 K and 7 T. This Faraday angle is much larger than the sensitivity of the measurements setup (about 2 mRad) and well above the measurement noise level, as shown in Figure 7. In addition to the 122 $cm^{-1}$  it is known that FePS3 has another excitation at higher energies of about 40 meV. Close examination of our Faraday angle rotation spectrum close to that energy reveals a very weak feature, which we attribute to this high-energy excitation. Figure 8 shows the Faraday rotation angle for selected temperatures above and below the antiferromagnetic transition of FePS3. Below $T_{N}$ but still close to the phase transition, it is very hard to detect any Faraday angle rotation above our measurement baseline and within the noise level however, as the system goes deep into the antiferromagnetic phase, a dip in the data appears at about 320 $cm^{-1}$ which corresponds to an energy of about 40 meV. The appearance of a Faraday rotation signal at $\leavevmode\nobreak\ 320$ $cm^{-1}$, combined with a temperature-dependent shift and intensity increase, provides strong evidence that this feature originates from a magnetic excitation.
Figure 8: Temperature evolution of the Faraday rotation angle in the 280–350  $cm^{-1}$ range of FePS3, measured at 7 T with FIR polarizer orientation Pol 2. A prominent feature near 320  $cm^{-1}$ emerges below the Néel temperature (TN = 118 K), increases in strength with decreasing temperature, and is attributed to a higher-energy mode at 40 meV. This temperature-dependent behavior further supports the magnetic origin of the excitation.
It is important to note here that the frequency 122 $cm^{-1}$ in our data corresponds to the energy of 15 meV detected at the commensurate $q=(0,1,0)$ momentum vector via inelastic neutron scattering measurements [17] with a momentum dispersion reaching about 17 meV at $q=(0,0.5,0)$. On the other hand, the excitation at about 40 meV is seen at the incommensurate momentum of, e.g., $q=(\pm 1,k,\mp 1/3)$ with a momentum dispersion reaching about 17 meV at e.g., $q=(\pm 2,k,\mp 2/3)$ (where $k$ is $0.5,1,1.5$). In general, optical spectroscopy is sensitive only to $q\approx 0$ collective excitations. Nevertheless, the frequencies at which we measure the Faraday angle rotation signal agree with the excitation energies of FePS3 as measured by inelastic neutron scattering. Moreover, it is possible that the difference in intensity between the signal at 122 $cm^{-1}$ (15 meV) and 320 $cm^{-1}$ (40 meV) is associated with the fact that the excitations are commensurate and incommensurate, respectively.
IV Discussion
IV.1 Electronic structure of FePS3
Figure 9: Local crystal field environment of Fe2+ in FePS3 and orientation of the $e_{g}$ (top) and $t_{2g}$ (bottom) orbitals. The $e_{g}$ states have one protuberance pointing toward the negative sulfur ions and one away from it. The $t_{2g}$ states point away from the negative ions. If we call the axis connecting the farthest corners of the two cubes the $z$ axis, the yzyz and zxzx orbitals have one protuberance along $z$ and one protuberance in the plane. The xyxy orbital has both protuberances pointing in the plane (i.e. away from the negative sulfur ions).
Fe2+, has a $3d^{6}$ configuration. The 210 possible $3d^{6}$ states are described by strongly correlated many-electron wavefunctions as a result of the Coulomb interaction. Thanks to the symmetries of the problem, these states can be labeled according to spin, orbital momentum, and total (spin plus orbital) angular momentum of the interacting electron system. In the absence of a crystal field and spin-orbit interaction, the ground state is described by the $^{5}D$ multiplet comprising $5\times 5$ degenerate eigenstates. The remaining 185 eigenstates are grouped in multiplets scattered over a range of 10 eV above the ground state [47].
In FePS3 the negatively charged sulfur ions surrounding Fe2+ occupy roughly the 8 corners of an octahedron (see Fig. 9). The ensuing crystal field splits the $^{5}D$ multiplet in $^{5}T_{2g}$ and $^{5}E_{g}$. The $^{5}E_{g}$ states have protuberances closer to the negative sulfur ions than the $^{5}T_{2g}$. Consequently, the energy of the $^{5}T_{2g}$ states is about 1 eV lower than that of the $^{5}E_{g}$ states.
Spin-orbit coupling splits the $^{5}T_{2g}$ states in three multiplets (see Fig. 10): The ground state has $J=1$, an excited $J=3$ multiplet at $\sim 30$ meV and $J=5$ at $\sim 60$ meV. A further refinement comes from the trigonal crystal structure, namely that two of the $^{5}T_{2g}$ states have one protuberance in the plane (i.e. away from the negative sulfur ions) and one perpendicular to it. The remaining $^{5}T_{2g}$ state has both protuberances pointing in the plane (i.e. away from the negative sulfur ions). The trigonal field $\Delta$ is then negative and separates the $^{5}T_{2g}$ multiplet into a $^{5}A_{g}$ ground state and a $^{5}E_{g}$ excited state. Refs. [23, 48, 49, 13, 50, 51, 52] also obtained/assumed $\Delta<0$. On the other hand, a positive sign was obtained in Ref. [53] by fitting high-temperature susceptibility data, see Fig. 10.
Figure 10: Tanabe-Sugano diagram of the $d^{6}$ energy levels in a trigonal crystal field adapted from Ref. [53]. The black arrow on the right is from the original paper. The green arrow on the left is our tentative assignment.
Although the orbital momentum is quenched by the crystal field, a finite orbital moment is induced by the spin-orbit coupling. This is illustrated in Fig. 10, which was adapted from Ref. [53]. The arrow on the right-hand panel is from the original paper, but in view of the expected sign of the crystal field discussed above, this seems unphysical. With this sign of crystal field the ground state of an isolated Fe2+ ion has no magnetic moment, whereas there is a moment for $\Delta<0$. For $\Delta>0$, the possibility remains that a magnetic moment is induced by the Weiss field [53].
The energy levels of our tentative assignment (green arrow in Fig. 10, $\Delta<0)$ are reproduced in Table 2. The highest energy states are consistent with observed absorption features in the range of 1000 cm-1. It is important to strike a note of caution at this point: the model used to generate this table did not take into account the orbital mixing of Fe $3d$ and the ligand states [54], and this cannot be remedied by a rescaling of interaction and/or crystal field parameters because the degree of orbital mixing will in general be different from one multiplet to another according to its shape and orientation. Consequently, the energy spacings in this Table 2 cannot be fully accurate.
Energy (meV) Degeneracy
126.6 2
126.5 1
120.2 2
61.4 2
41.4 2
26.9 2
17.9 1
6.4 1
0 2
Table 2: Energies of the $^{5}T_{2g}-$manifold split by crystal field and spin-orbit parameter, corresponding to the yellow arrow in the left panel of Fig. 10
The ground state is a spin-orbit doublet with a finite magnetic moment. In an AF environment the exchange field of the surrounding ions lifts the degeneracy of the ground state, making the spin point up or down. Applying a magnetic field along $z$ has the effect for half of the spins that their energy increases, and for the other half it decreases. If one now observes the transition from the ground state to the singly degenerate excited state at 17.9 meV, the spectrum will show a splitting of this transition, the peak moving up coming from the down-spin atoms, and the peak moving down coming from the up-spin atoms.
Another potential interesting feature is the transition from the ground state to the singly degenerate 6.4 meV excited state. Formally this can be identified with the anti-ferromagnetic resonance [55]. In Ref.  [55] the phenomenon is described as the precession of the spin in an anisotropy field $H_{a}$ with frequency $\hbar\omega/\gamma=H_{0}\pm[H_{a}(2H_{e}+H_{a})]^{1/2}$ where $H_{0}$ is the external field, $\gamma=ge/2mc$ and $H_{e}$ is the Weiss exchange field. In the local impurity case $H_{a}$ is replaced by the splitting of the three-fold degenerate spin-orbital state in a ground state doublet and an excited singlet state. It is important to emphasize that what we usually call the “magnon” in an optical spectrum is the lowest excitation, in the present case, 6.4 meV. The state at 17.9 meV constitutes a rearrangement of the many-body wave-function compared to the ground state, and is best referred to as a $dd$-exciton.
IV.2 Crystal Symmetries in FePS3
We previously showed in Fig. 3 a comparison between our polarized transmission data and Raman microscopy done on similar FePS3 of the same growth batch. Several vibrational modes of FePS3 exhibit a clear Faraday angle rotation, providing strong evidence for spin–phonon coupling. This observation is consistent with earlier reports [37]. Interestingly, the Raman-active phonon mode at 108 $cm^{-1}$ appears as a weak absorption feature in the infrared magneto-transmission data (See Fig.  3 and Table 1), where it also develops a measurable Faraday rotation of approximately 25 mrad. The same mode is also showing an avoided crossing behavior with the 122 $cm^{-1}$ excitation at high magnetic fields [37].
The appearance of the Raman modes in the IR data suggests a breaking of inversion symmetry and the mixing of odd and even collective modes. The application of a magnetic field naturally breaks time-reversal symmetry and can result in the appearance of Raman excitations in our magneto-transmission measurements. This observation is also shown in the high field magneto transmission data of Ref.  [56], particularly the avoided crossing of the 122 $cm^{-1}$ excitation and the Raman mode in the IR data. However, there is no conventional interpretation of why this specific Raman mode will also develop a Faraday angle rotation.
Two additional modes show a different behavior, which also cannot be attributed to magnetic field-induced time-reversal symmetry (TRS) breaking. The modes at 133 $cm^{-1}$ and 166 $cm^{-1}$ are infrared active modes (see comparison with the Raman data in Fig. 3 and Table 1) but develop a Faraday angle rotation of about 25 mRad and 50 mrad, respectively, at 7 T as seen in Fig. 7. Therefore, these modes already show a clear spin-phonon coupling through the Faraday angle rotation at low magnetic fields. It is possible that the phonon symmetry in this case is coupled to the in-plane anisotropy of the magnetic order, being different along the $a$-axis and along the $b$-axis, resulting in magnetic field-induced chirality of the phonons’ optical response [57].
Figure 11 shows a color map of the polarized mid-infrared transmission data in our FePS3 single crystals and another intriguing spectral feature which is centered around 900 $cm^{-1}$. A wide absorption, and also a wide peak in the optical conductivity, can be seen mainly in the Pol2 data and below the AFM phase transition. Similarly, a peak in the Raman spectroscopy data of our samples is observed below the AFM phase transition and around a Raman shift of about 900 $cm^{-1}$.
Figure 11: Top panel: Color map of polarized infrared transmission (from 50  $cm^{-1}$and for Pol 1) at zero field as a function of temperature. Bottom panel: 2D color map of the unpolarized Raman Intensity in zero field as a function of temperature. The temperature range in both measurements was obtained between 10 K to 150 K in intervals of 10 K.
Therefore, the appearance of the 122 $cm^{-1}$ excitation and the mid-infrared excitation at 900 $cm^{-1}$, all visible in both infrared and Raman spectra at zero field, suggests that inversion symmetry is broken in FePS3. This claim is in line with recent experimental indications [58] and in contrast to second harmonic generation measurements [39], which argue that FePS3 preserves inversion symmetry in the magnetic state.
IV.3 Dichroism of the Optical Circular Conductivity
FePS3 is considered to have an Ising-like (1D) spin Hamiltonian where the spins are directed along the $c$-axis. Hence, there should be a negligible, if any, anisotropy effect on the absorption spectra in different polarizations along the crystal’s $ab$ plane, as can be seen in our data (see App. A). This allows us to consider the 122 $cm^{-1}$ as an isotropic excitation in the polarized magneto-transmission data. Since we can quantitatively obtain the absolute transmission, the complex conductivity, and the Faraday angle of FePS3 in our measurements, it is, therefore, possible to extract from our data the circular optical conductivity of the split 122 $cm^{-1}$ mode. The two field-separated collective modes can be associated with the right- and left-handed circular polarizations of the two opposite spin sublattices in the antiferromagnetic state  [59]. At zero field, there is no difference between one spin population and the other, and the 122 $cm^{-1}$ excitation has a single eigen energy. Hence, the two circular polarizations combine into one linearly polarized mode at 122 $cm^{-1}$. The application of a magnetic field splits the modes into low and high energy modes with a corresponding circular optical conductivity $\sigma_{+}$ and $\sigma_{-}$.
We follow the procedure described in Ref. [22] however, due to the anisotropy in the magnetic state, the in-plane optical conductivity tensor should be described as follows:
$\sigma=\begin{pmatrix}\sigma_{xx}&\sigma_{xy}\\ -\sigma_{xy}&\sigma_{yy}\end{pmatrix}$ (2)
where the in-plane anisotropy of the FePS3 optical conductivity is given by different dynamic conductivity functions $\sigma_{xx}$ and $\sigma_{yy}$. The latter terms capture the ordinary absorption and the birefringence of the sample, while the diagonal term, $\sigma_{xy}$, is associated with the magneto-optical effects and the optical Hall conductivity. Plugging in the eigenvectors of the Right- and Left-handed circular polarized light $e_{\pm}=1/\sqrt{2}\left(1,\pm i\right)$ results in the circular optical conductivity:
$\sigma_{\pm}(\omega)=\frac{\sigma_{xx}(\omega)+\sigma_{yy}(\omega)}{2}\pm i\sigma_{xy}(\omega)$ (3)
for the Right-handed ($\sigma_{+}$) and Left-handed ($\sigma_{-}$) optical circular conductivity. It is important to note here that we work under the assumption that the monoclinic tilt is non-negligible and therefore take the orthorhombic approximation because the optic axes in our measurement stay nearly in-plane. We also note that $\sigma_{xy}$ should be independent of the actual axis in which we have resolved the Faraday angle rotation. Moreover, while linear birefringence mixes the two circular eigenmodes, so they are not perfectly circular, it does not affect the Faraday angle rotation since it is not related to the off-diagonal terms and cancels out in the cross-polarization measurement. However, the values of $\sigma_{xy}$ can be different for the measurement and analysis of the Faraday angle rotation in Pol1 vs Pol2 due to systematic errors, e.g. residual cross polarization signal, and/or crystal misalignment. Our results show that throughout most of the spectral range, including the lower energy mode (approx. 116 $cm^{-1}$) and phonons, there is a perfect agreement between the two sets of data with respect to $\sigma_{xy}$ obtained from the two in-plane polarizations. Some discrepancy between the values can be detected around the upper energy mode (approx. 129  $cm^{-1}$, see App. B), which shows that it is local in frequency and probably not related to systematic errors or any sample tilt. Therefore, we have averaged the two $\sigma_{xy}$ outputs in order to obtain the optical circular conductivity $\sigma_{\pm}$.
Figure 12 shows the circular optical conductivity obtained from the analysis of the Magneto-Optical Kerr/Faraday Effect [36] and by taking into account the in-plane optical conductivity tensor as described above. As expected from a collinear antiferromagnetic resonance, there is a clear dichroism between $\sigma_{+}$ and $\sigma_{-}$ under the application of a magnetic field seen at frequencies of about 116 $cm^{-1}$ and 129 $cm^{-1}$ for 7 T for the lower energy and upper energy, respectively, of the 122 $cm^{-1}$ zero-field mode.
In fact, Fig.  12 shows that the lower mode is completely dichroic since the spectral weight of $\sigma_{-}$ in that excitation is the same magnitude as that of $\sigma_{1}$ at zero field while $\sigma_{+}$ shows a clear peak. On the other hand, the upper energy mode at around 129 $cm^{-1}$ does not show an ideal dichroic response with $\sigma_{+}$ being about factor of 1/3 from the value of $\sigma_{-}$. This response is not subject to the average of $\sigma_{xy}$ of the two polarizations and can be seen in both sets of data for the upper energy mode (129 $cm^{-1}$).
Since this anomalous dichroism appears only in the upper energy mode (129  $cm^{-1}$) which is also close to an infra-red active phonon, we suggest that spin-phonon coupling is responsible for prohibiting optical transitions between the two circular energy states, thus giving a residual, but still significant, spectral weight to $\sigma_{+}$ which is supposed to be closer to $\sigma_{1}$ at zero field in that area (around $1\leavevmode\nobreak\ S$  $cm^{-1}$). Previous reports of coupling between the lower energy mode (116  $cm^{-1}$) and an optical phonon have been reported as an avoided crossing in the optical spectra [56]. Therefore, we suggest that an avoided transition due to spin-phonon coupling involving the upper energy mode (129  $cm^{-1}$) and the infrared active phonon at 133  $cm^{-1}$is already present at low fields.
Figure 12: Circular optical conductivity $\sigma_{+}$ and $\sigma_{-}$ of FePS3 at 5 K and 7 T, obtained by averaging the off-diagonal conductivity $\sigma_{xy}$ extracted independently for Pol 1 and Pol 2 using Eq. 3.
A careful look into $\sigma_{\pm}$ exposes several other excitations which develop a dichroic response. Since the magnetic field is breaking the time-reversal symmetry, it is also possible to measure the circular dichroism of the Raman active mode at 108 $cm^{-1}$, which appears in our far infrared spectroscopy data and previous magneto-transmission measurements [56]. Figure 12 shows that the 108 $cm^{-1}$mode develops a Faraday angle rotation, which is above our measurement sensitivity. Another Raman active mode at 160 $cm^{-1}$ can be detected in the circular optical conductivity in both polarizations due to a possible inversion symmetry breaking and leakage of the even vibrational modes into the odd modes IR spectra. Since the 160 $cm^{-1}$mode is close to one of the IR active modes in our measurement, the Faraday angle of this mode is seen better in the second principal axis polarization shown in the right panel of Fig. 7. In both cases, our data allow us to extract the spectral weight of the circular conductivity of these modes.
V Conclusions
We reported a comprehensive magneto-optical study of FePS3 that resolves the interplay between magnetic and lattice degrees of freedom. In the antiferromagnetic state, the phonon spectrum becomes strongly anisotropic. On the other hand, the $122$  $cm^{-1}$ ($15$ meV) excitation is polarization-independent, hardens with an order-parameter-like temperature dependence, and splits with field at low temperatures, producing a large Faraday rotation. By combining absolute transmission with full-protocol Faraday measurements, we extracted the circular optical conductivities and uncovered a clear dichroism of the split excitations. The asymmetric spectral weight of the upper branch near $129$  $cm^{-1}$ is naturally explained by spin–phonon coupling to a nearby infrared phonon. Sizable Faraday rotation on several phonons, together with the appearance in infrared of Raman-active modes and a broad band near $900$  $cm^{-1}$ below $T_{\mathrm{N}}$, indicates inversion-symmetry breaking in the magnetic phase. A weak Faraday signature around $320$  $cm^{-1}$ ($\sim 40$ meV) points to an additional magnetic excitation. These findings highlight a rich magneto-optical landscape in FePS3, motivate future experiments combining ultrafast optical probes or monolayer extensions, and demonstrate a general route to quantify spin–lattice coupling and circular dichroism in two-dimensional antiferromagnets.
Acknowledgments
This research was supported by the Israel Science Foundation (ISF) through project number 666/23. N.B. would like to acknowledge useful discussions with J. Teyssier, G. Blumberg, and M. Orlita.
Appendix A Transmission Data
As explained in the main text and in Sec. II, FePS3h̃as an anisotropic response in the antiferromagnetic state. In order to identify two principal optical axes for our polarization-dependent measurements, we have measured the transmission spectrum of our sample at 5 K (deep in the AFM state) and at zero field as a function of the incoming polarizer angle. Figure 13 shows a colormap plot of the polarization-dependent transmission spectra in which two distinct spectra are identified as the principal optical axes of the FePS3s̃ingle crystal sample. The same procedure was conducted in the MIR range using a KRS-5 polarizer and associated with their counterpart FIR polarized data. The entire set of transmission data from the FIR and MIR ranges was then combined into two sets termed as "Pol 1" and "Pol 2" in our work.
Figure 13: Colormap plot of the transmission spectra as a function of the incoming polarizer (Gold polarizer for the FIR range) angle in the range of 50 to 250  $cm^{-1}$. The gray dashed lines represent the two principal polarizations $P$ used in this work for the temperature and magnetic field-dependent transmission measurements. Here $P=150^{\circ}$ is defined as "Pol 1" and $P=60^{\circ}$ as "Pol 2". The 122 $cm^{-1}$ excitation can be seen in this colormap as a polarization-independent absorption compared to the lattice collective modes.
In Figure 14 (a)-(b) we show a comparison of the transmission data measured for the two principal polarizations at 150 K, which is above the Néel temperature, TN=118 K, and at 5 K, which is deep in the antiferromagnetic state. At 150 K, there is no significant difference between the two spectra, except for the phonon at around 220 $cm^{-1}$, indicating isotropic transmission behavior in the paramagnetic phase and on the aba-b plane of the monoclinic structure of FePS3. However, as the temperature approaches TN, noticeable changes appear in the spectra, making it impossible to fit the data with a single isotropic model and therefore pointing to the onset of the magnetic order and increasing fluctuations of the magneto-elastic effects. The same spin fluctuations are also responsible for the change in the mode around 100 $cm^{-1}$ seen in the Raman spectrum [43]. Below $T_{N}$, there is an abrupt anisotropy in the spectra up to about 250  $cm^{-1}$ since the three-fold rotational symmetry in FePS3 is broken by the formation of the zigzag spin chains, resulting also in the appearance of Brillouin zone folded phonons as was seen also previously in Raman spectroscopy [15, 14].
To emphasize the effect of the magnetic state on the anisotopic optical response of FePS3  we present in Figure 14(c) the zero-field transmission data for MnPS3 at two temperatures: 100 K ($T>T_{N}$) and 10 K ($T\ll T_{N}$), where $T_{N}=78$ K. The spectra show that in the antiferromagnetic state of MnPS3, there are no new phonon modes. Both MnPS3 and FePS3 have the same monoclinic crystal structure with a C2/m space group. However, their magnetic structure in the antiferromagnetic state is different. In FePS3, the unit cell doubles along one direction due to the zigzag magnetic structure and the formation of ferromagnetic chains along the $a$-axis and antiferromagnetic coupling between the chains along the $b$-axis. As a result, new phonon modes, which are folded from higher momentum points to the $\Gamma$ point, emerge below $T_{N}$ [15, 14, 60, 43]. In contrast, MnPS3 has an isotropic Néel spin configuration, meaning that each magnetic moment is antiparllel to all its nearest neighbors. As a result, there is no need to consider a larger unit cell in order to incorporate this structure, as happens in FePS3. Therefore, no new phonon modes appear in its spectrum, consistent also with the Raman spectroscopy measurement [61].
Figure 14: Transmission spectra of FePS3 ($T_{N}=118$ K) for two polarizations at (a) 150 K ($T>T_{N}$) and (b) 5 K ($T\ll T_{N}$). (c) Zero-field transmission spectra at 100 K ($T>T_{N}$) and 10 K ($T\ll T_{N}$) for MnPS3 ($T_{N}=78$ K). No additional phonon modes emerge in the antiferromagnetic state.
Appendix B Optical Circular Conductivity in the Birefringence Case
We start by defining a dielectric tensor of an anisotropic sample including the off-diagonal magneto optical terms which is given by:
$\epsilon(\omega)=\begin{pmatrix}\epsilon_{xx}&\epsilon_{xy}&0\\ \epsilon_{yx}&\epsilon_{yy}&0\\ 0&0&\epsilon_{zz}\end{pmatrix}$ (4)
where $x,y,z$ are the cartesian axes. Since the light is propagating in the $z$ direction, we can refer to the in-plane dielectric tensor $e_{\perp}$, which is a $2\times 2$ matrix, in the following derivation. It is important to note that our derivation follows the orthorhombic approximation since we neglect, at the moment, the tilt of the monoclinic structure of FePS3. Therefore, the dielectric tensor in the xyx-y plane:
$\epsilon_{\perp}=\begin{pmatrix}\epsilon_{xx}&\epsilon_{xy}\\ -\epsilon_{xy}&\epsilon_{yy}\end{pmatrix}$ (5)
where the in-plane anisotropy is given by the fact that $\epsilon_{xx}\neq\epsilon_{yy}$ and the off diagonal element is $\epsilon_{yx}=-\epsilon_{xy}$.
In the same way we can define the in-plane optical conductivity tensor:
$\sigma{\perp}=\begin{pmatrix}\sigma_{xx}&\sigma_{xy}\\ -\sigma_{xy}&\sigma_{yy}\end{pmatrix}$ (6)
since the dielectric function and optical conductivity are connected via:
$\epsilon_{\mu\nu}(\omega)=\delta_{\mu\nu}+\frac{i}{\epsilon_{0}\omega}\sigma_{\mu\nu}(\omega)$ (7)
where $\mu,\nu\in{x,y}$, $\delta_{\mu\nu}$ is the Kronecker delta, $\epsilon_{0}$ is the vacuum permittivity, and $\omega$ is the optical frequency.
The circular polarization basis is given by the eigenvectors:
$\hat{e}_{R,L}=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ \mp i\end{pmatrix}$ (8)
where $L$ represents a clockwise rotation or Left-Handed (LH) circular polarization, while $R$ represents a counterclockwise rotation or Right-Handed (RH) circular polarization.
In order to convert from the xyx-y basis to the circular basis, we need to find the matrix $U$ which will convert it as follows:
$\begin{pmatrix}\hat{x}\\ \hat{y}\end{pmatrix}=U\begin{pmatrix}\hat{e}_{R}\\ \hat{e}_{L}\end{pmatrix}$ (9)
which is:
$U=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ i&-i\end{pmatrix}$ (10)
while for the inverse case:
$\begin{pmatrix}\hat{e}_{R}\\ \hat{e}_{L}\end{pmatrix}=U^{\dagger}\begin{pmatrix}\hat{x}\\ \hat{y}\end{pmatrix}$ (11)
we write that:
$U^{\dagger}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-i\\ 1&i\end{pmatrix}$ (12)
where $U$ is a unitary matrix and $U^{\dagger}U=UU^{\dagger}=\mathbb{1}$
We can now convert the dielectric tensor from the $xy$ basis to the circular basis as follows:
$\epsilon_{circ}=U^{\dagger}\epsilon_{\perp}U=\begin{pmatrix}\epsilon_{+}&\delta\\ \delta&\epsilon_{-}\end{pmatrix}$ (13)
where the diagonal elements are:
$\epsilon_{\pm}=\frac{\epsilon_{xx}+\epsilon_{yy}}{2}\pm i\epsilon_{xy}$ (14)
and the off-diagonal elements, representing the birefringence, are:
$\delta=\frac{\epsilon_{xx}-\epsilon_{yy}}{2}$ (15)
Here the (++) sign refers to the LH circular dielectric element, $\epsilon_{+}=\epsilon_{LL}=\hat{e}^{\dagger}_{L}\epsilon\hat{e}_{L}$, the (-) sign refers to the RH circular dielectric element, $\epsilon_{-}=\epsilon_{RR}=\hat{e}^{\dagger}_{R}\epsilon\hat{e}_{R}$, and $\delta=\epsilon_{RL}=\epsilon_{LR}$.
If $\delta=0$ or $\epsilon_{xx}=\epsilon_{yy}$ then the eigenmodes of the dielectric tensor are purely circular where $n_{L}^{2}=\epsilon_{+}=\epsilon_{LL}$ for LH circular polarization and $n_{R}^{2}=\epsilon_{-}=\epsilon_{RR}$ for RH circular polarization and $n=\sqrt{\epsilon}$ is the refractive index. Ref. [22] follows this isotropic case in order to extract the Faraday (Kerr) angle rotation from the magneto-transmission (-reflection) data. However, since FePS3 is anisotropic, we have to go through the entire derivation of the Faraday angle in the case of a birefringent sample.
For linear polarized light which is a superposition of LH and RH circular polarization, we can write that:
$\vec{E}_{in}=E_{0}\hat{x}=\frac{E_{0}}{\sqrt{2}}(\hat{\epsilon}_{L}+\hat{\epsilon}_{R})$ (16)
The incoming field propagates through a slab of thickness $d$, where each circular component will accumulate a different phase:
$\vec{E}_{out}=\frac{E_{0}}{\sqrt{2}}(\hat{\epsilon}_{L}e^{ikn_{L}d}+\hat{\epsilon}_{R}e^{ikn_{R}d})$ (17)
The phase difference $\Delta\phi=\frac{\omega d}{c}(n_{L}-n_{R})$ between the two modes will result in the Faraday rotation angle:
$\theta_{F}=\frac{1}{2}\Re{\left[\frac{\omega d}{c}(n_{L}-n_{R})\right]}$ (18)
and the ellipticity:
$\eta_{F}=\frac{1}{2}\Im{\left[\frac{\omega d}{c}(n_{L}-n_{R})\right]}$ (19)
A nonzero $\delta$ implies that the sample has a linear birefringence term and the dielectric tensor in the circular basis is not diagonal, nor $\hat{\epsilon}_{R}$ and $\hat{\epsilon}_{L}$ are the true purely circular eigenmodes. The polarization eigenstates in this case tilt away from the circular axes and mix the LH and RH components.
The eigenvalues of $\epsilon_{circ}$ are therefore obtained by diagonalizing the matrix $\mathrm{det}(\epsilon_{circ}-\lambda_{\pm}\mathbb{1})=0$ but recognize that $n_{\pm}=\sqrt{\lambda_{\pm}}$, Therefore we can write the refractive indices as::
$n^{2}_{\pm}$ $=\frac{\epsilon_{+}+\epsilon_{-}}{2}\pm\sqrt{\left(\frac{\epsilon_{+}-\epsilon_{-}}{2}\right)^{2}+\delta^{2}}$ (20)
$=\frac{\epsilon_{xx}+\epsilon_{yy}}{2}\pm\sqrt{\delta^{2}-\epsilon_{xy}^{2}}$ (21)
where the first term in the right-hand side of Eq. 21 is the average of the in-plane dielectric functions:
$\bar{\epsilon}=\frac{\epsilon_{xx}+\epsilon_{yy}}{2}$ (22)
Below the square root are the linear birefringence term:
$\delta=\frac{\epsilon_{xx}-\epsilon_{yy}}{2}$ (23)
and the off-diagonal term which corresponds to the Magneto-Optical effect on the optical functions of the sample. Therefore we can write Eq. 21 as follows:
$n^{2}_{\pm}=\bar{\epsilon}\pm\sqrt{\delta^{2}-\epsilon_{xy}^{2}}$ (24)
We can define $\epsilon_{xx}=n_{0}^{2}e^{+\theta}$ and $\epsilon_{yy}=n_{0}^{2}e^{-\theta}$ with a common scale $n_{0}^{2}=\sqrt{\epsilon_{xx}\epsilon_{yy}}$ and an exponential factor $\theta=0.5\ln(\epsilon_{xx}/\epsilon_{yy})$ in order to extract $n_{\pm}$. In fact, $n_{0}$ is the average refractive index or the geometric mean of $n_{x}=\sqrt{\epsilon_{xx}}$ and $n_{y}=\sqrt{\epsilon_{yy}}$, and $\theta$ represents the in-plane anisotropy, with $\theta=0$ being the isotropic case. We can then express in a different way the average of the in-plane dielectric functions as:
$\bar{\epsilon}=n_{0}^{2}\cosh{\theta}$ (25)
and the linear birefringence term as:
$\delta=n_{0}^{2}\sinh{\theta}$ (26)
and write Eq. 24 as:
$n^{2}_{\pm}=n_{0}^{2}\left(\cosh{\theta}\pm\sqrt{\sinh^{2}{\theta}-\alpha^{2}}\right)$ (27)
with $\alpha=\epsilon_{xy}/n_{0}^{2}$ being a dimensionless antisymmetric parameter and much smaller than 1 for typical magneto-optical experiments.
For the isotropic case (as in Ref. [22]) where $\epsilon_{xx}=\epsilon_{yy}=\epsilon$ so that $\delta=0$ Eq. 24 can be simplified to $n^{2}_{\pm}=\epsilon\pm\epsilon_{xy}$ and expanded to obtain:
$n_{\pm}=n_{0}\pm\frac{\epsilon_{xy}}{2n_{0}}+\mathcal{O}(\epsilon_{xy}^{2})$ (28)
with a linear in $\epsilon_{xy}$ term. The eigenmodes acquire a phase:
$\phi_{\pm}=k_{\pm}d=\frac{\omega n_{\pm}d}{c}$ (29)
through a slab of thickness $d$ where $c$ is the speed of light in vacuum and $k_{\pm}$ is the wavenumber for RH and LH circular polarization. Plugging in Eq. 28 and $\phi_{0}=\omega n_{0}d/c$ gives:
$\phi_{\pm}=\phi_{0}\pm\frac{\omega d}{2c}\frac{\epsilon_{xy}}{n_{0}}$ (30)
The complex Faraday angle is defined as:
$\Theta=\theta_{F}+i\eta=\frac{1}{2}\left(\phi_{+}-\phi_{-}\right)$ (31)
and therefore allow us to write the relation between the off-diagonal terms of the dielectric tensor to the Faraday angle:
$\epsilon_{xy}(\omega)=\frac{2c}{\omega d}n_{0}(\omega)\Theta(\omega)$ (32)
or the same for the conductivity tensor via:
$\sigma_{xy}(\omega)=-i\omega\varepsilon_{0}\epsilon_{xy}(\omega)$ (33)
This derivation highlights the direct relation between the linear in $\epsilon_{xy}$ term to the Faraday angle rotation and the fact that the off-diagonal terms are those responsible for it.
If the crystal is slightly birefringent so that $\delta\ll\bar{\epsilon}$ the square root in Eq. 24 can be expanded as:
$\sqrt{\delta^{2}-\epsilon_{xy}^{2}}=\delta-\frac{\epsilon_{xy}^{2}}{2\delta}+\mathcal{O}(\epsilon_{xy}^{4})$ (34)
which does not have any linear in $\epsilon_{xy}$ term. The Faraday angle rotation is obtained from measurements in positive and negative magnetic fields. Therefore, the contribution from the birefringent term will be canceled out since (i) the magnetic field does not affect $\delta$ directly, (ii) $\epsilon_{xy}$ is squared and symmetric for negative and positive fields. As a result, the birefringent term will not contribute to the total Faraday angle rotation. In particular, if $\delta\ll\epsilon_{xy}$ then the rotation is still dominated by $\epsilon_{xy}$ and the circular basis is a good approximation. However, if $\delta\gg\epsilon_{xy}$ linear birefringence dominates and the Faraday effect is strongly suppressed or masked, which is not our case, as can be seen in the data.
Figure 15: Top panel: Circular optical conductivities $\sigma_{+}$ and $\sigma_{-}$ reconstructed separately for each polarization (Pol 1 and Pol 2), using the combination of full-protocol Faraday rotation and absolute transmission spectra fitted using RefFit software. Bottom panel: Circular optical conductivities $\sigma_{+}$ and $\sigma_{-}$ reconstructed for each polarization (Pol 1 and Pol 2) using Eq. 3, where the Hall conductivity $\sigma_{xy}$ was independently derived for each polarization without averaging.
To further demonstrate the negligible effect of the birefringence on the Faraday rotation, we refer to our analysis of the polarized data. The main text considers an average of the dielectric constant in order to obtain the optical circular conductivity from the Faraday angle rotation and magneto transmission measurements in the two principal axes. Figure 15 provides complementary representations of the circular optical conductivity, prior to the averaging process, which help assess the consistency of the analysis presented in the main text (Fig. 12).
Figure 15 (top) shows $\sigma_{+}$ and $\sigma_{-}$ reconstructed separately for the two principal polarizations (noted by the two angles of the analyzer). In this case, the circular conductivities were obtained by combining the Faraday rotation measured using the full rotation protocol with the absolute transmission spectra, and fitting them directly using the MOKE + VDF procedure in RefFit software.
Figure 15 (bottom) displays $\sigma_{+}$ and $\sigma_{-}$ for the two polarizations however now derived by using Eq. 3 where $\sigma_{xx}$ and $\sigma_{yy}$ were averaged. Here, the off-diagonal term $\sigma_{xy}$ was determined independently for each analyzer configuration without averaging.
Fig. 12 of the main text shows the result obtained by averaging $\sigma_{xy}$ from the two polarizations, providing a single optical circular conductivity spectrum for FePS3. Comparing the different plots in Fig. 15 demonstrates that the overall important features of the circular dichroism that we noted in the main text, including the splitting of the 122  $cm^{-1}$ mode and the spectral weight asymmetry, are reproducible across all different analysis procedures. Minor differences between the reconstructions mainly reflect systematic effects such as residual cross-polarization signals and slight misalignment of the sample.